claims are uniformly distributed over the range(0,10).if four claims are selected at random, determine the probability that the second largest claim is less than 6
Answer #1

Solution :

Given that,

a = 0

b = 10

P(x < c) = (c - a) / (b - a)

P(x < 6) = (6 - 0) / (10 - 0)

= 6 / 10 = 0.6

Probability = 0.6